OpenStudy (anonymous):

f(x)= |4-x|-|x+3|+3x-5

OpenStudy (anonymous):

and the question is?

OpenStudy (anonymous):

rewrite without absolute values

OpenStudy (anonymous):

f(x)=(4-x)-(x+3)+3x-5 f(x)=(4+x)-(-x+3)+3x-5 then you just simplify both by putting like terms together

OpenStudy (anonymous):

I'm confused, how.... why? u just switch the xs?

OpenStudy (anonymous):

yep for example when you have |x| could be - or +

OpenStudy (anonymous):

so you have to do both versions

OpenStudy (anonymous):

what about the normal #s, like 3 and 4, you didnt negate them?

OpenStudy (anonymous):

did i not put one? Dx

OpenStudy (anonymous):

what's going on now?

OpenStudy (anonymous):

hello?